OpenStudy (anonymous):

compute the derivative of y=ln(3+sin(x^2))

OpenStudy (anonymous):

\[= \frac{ \frac{d}{dx} (3+\sin(x^2) ) }{ 3+\sin(x^2)}\]

OpenStudy (anonymous):

\[= \frac{ 2x \cos(x^2)}{3+\sin(x^2)}\]

OpenStudy (anonymous):

if \[ y= \ln [ f(x)] \rightarrow \frac{dy}{dx} = \frac { f ' (x)}{f(x)}\]